Subject: Thinfilms morning report 11/12/02
From: Jim Haydon <jhaydon@snf.stanford.edu>
Date: Tue, 12 Nov 2002 09:20:47 -0800




Balzers:    ok
Gryphon:    Ion gauge controller  is not working, replacement may be
available as early as tomorrow 11/13/02
Implanter:    ok
Innotec:     ok
Metalica:    I would like regen the Cryo pump later today over night so
it may not be ready until late Wed morning 11/13/02
STS Dep:    Seal replacement Wed afternoon 11/13/02
wbgen- CTB:    ok
wbgen-LHP:    ok
wbgen-RHP:    ok
wbMETAL:    ok
wbNITRIDE:    ok
wbNONMETAL:    ok
wbDIFF:    ok
wbSILICIDE:    ok
wbGAASCL:    ok
wbMISCRES:    ok
wbOPTRES:    ok
DI WATER:    ok
NITROGEN:    ok